INFRARED SPECTRUM OF THE PRODUCTS OF THE INTERACTION OF EXCITED NEON ATOMS WITH CCl_{4} TRAPPED IN SOLID NEON

CATHERINE L. LUGEZ AND \underline{MARILYN E. JACOX}, National Institute of Standards and Technology, Gaithersburg, MD 20899.

When a Ne:CCl_{4} sample is codeposited at approximately 5 K with a beam of neon atoms that have been excited in a microwave discharge, the infrared spectrum of the resulting solid deposit shows prominent absorptions that have previously been assigned^{a} to Cl_{2}CCl–Cl, as well as other absorptions of uncharged, cationic CCl_{n} (n = 1-3) and anionic CCl_{n} (n = 3,4) species. Studies of Ne:^{1}3CCl_{4} samples and studies of the photodestruction of the products when the deposit is subjected to various bands of visible and ultraviolet radiation support the proposed assignments.

aG. Maier, H. P. Reisenauer, J. Hu, B. A. Hess, Jr., and L. J. Schaad, Tetrahed. Lett. 30, 4105 (1989).